Research Scenario #2:

A researcher wanted to study the effect of alcohol on reaction time. She hypothesized that alcohol will INCREASE reaction time (participants will take longer to react). She selected a sample of n= 36 participants from Rutgers University. The 36 participants each consumed a 6-ounce glass of wine. Thirty minutes later, the researcher measured each participant’s reaction time, using a standardized driving simulation task for which the regular population has a µ = 400 msec reaction time with a σ = 48. The reaction time mean for the sample was M= 412 msec. Are the data sufficient to conclude that the alcohol significantly increased reaction time?

Based on the above research scenario, please answer the following questions:

  1. What is the appropriate hypothesis test? __________________

  1. What two means are you comparing in this test? _____________________________

  1. Please calculate the appropriate hypothesis test using all four steps:

Step 1:

Step 2:

Step 3:

Step 4: _______________________________

Write the statistical statement for the results: __________________________________

Interpret your results (relating back to the hypothesis): _______________________________________________________________________________________________________________________________________________________________________________________________________________________________________

If appropriate, please compute effect size:   Answer: _______
Answer #1

Ans:

One sample mean test(z test)

We are comparing sample mean to a hypothised mean value of 400.

alpha=0.05

critical z value=1.645

Decision rule:Reject H0,if z>1.645

Test statistic:

z=(412-400)/(48/sqrt(36))

z=1.5

p-value=P(z>1.5)=0.0668

As,p-value>0.05,we fail to reject the null hypothesis.

There is not sufficient evidence to conclude that the alcohol significantly increased reaction time.

Effect size=(412-400)/48=0.25